# X-Vine Models for Multivariate Extremes

### v1

@@ 1 Introduction
For copula-based dependence modelling, outside the context of extreme value analysis, vine constructions have grown into a versatile and widely applied approach (Czado, 2019; Czado and Nagler, 2022; Nagler and Vatter, 2023). Our contribution is to make vine-based methods operational, in theory and practice, for densities of exponent measures of multivariate extreme value distributions. The challenge to overcome is that these densities do not integrate to one but to infinity. A change of margin transforms exponent measure densities into tail copula densities, which still have infinite mass, but with structural properties that resemble copula densities more closely.
@@
The models constructed in this way turn out to be the tail limits of regular vine copulas. For the special case of D-vine copulas, such tail limits were already computed in Joe (1996) and Joe et al. (2010), focusing on tail dependence functions rather than densities. In Simpson et al. (2021), tails of D-vine and C-vine copulas were investigated from the perspective of the limiting shapes of sample clouds.
X-vine constructions are related to but different from recently proposed graphical models for extremes, either for directed or undirected graphs (Gissibl and Klüppelberg, 2018; Engelke and Hitz, 2020; Engelke et al., 2022). Markov trees (Segers, 2020; Hu et al., 2023) are a special case in the intersection of graphical and X-vine models.
After reviewing background on multivariate extremes and regular vine sequences in Section 2, we state a version of Sklar’s theorem for tail copula densities in Section 3. Some well-known parametric examples are worked out in Section 4, with a focus on the simplifying assumption that the conditional copula densities only depend on the values of the conditioning variables through their indices, but not their actual values. Section 5 contains the paper’s main results, showing on the one hand how a general tail copula density can be decomposed along any regular vine sequence, and, on the other hand, how to construct a tail copula density from a regular vine sequence and bivariate building blocks. The so-called X-vine models arising in this way are put to work in the subsequent sections, covering simulation algorithms (Section 6), parameter estimation and model selection (Section 7), simulation studies (Section 8), and a case study on US flight delay data (Section 9) taken from Hentschel et al. (2022). Section 10 concludes. The appendices contain a detailed example to illustrate the theory of regular vine sequences and X-vine models, the proofs of the paper’s results, additional theory on Sklar’s theorem, exponent measure densities and the connection with regular vine copulas, as well as additional numerical results. The methods are implemented in R (R Core Team, 2023), relying in particular on packages graphicalExtremes (Engelke et al., 2022) and VineCopula (Nagler et al., 2023); a dedicated software package is under construction.
@@ 2.1 Multivariate extreme value theory: tail copulas and their densities
Tail copulas
Classical extreme value theory starts from the assumption that the distribution function $F$ of a random vector $\bm{X}=(X_{1},\ldots,X_{d})$ is in the max-domain of attraction of a multivariate extreme value distribution. This assumption comprises two parts: (i) all univariate marginal distributions $F_{1},\ldots,F_{d}$ are in the domains of attraction of some univariate extreme value distributions; (ii) after reduction of the univariate margins to a common scale, the joint distribution of the random vector is multivariate regularly varying. We refer to Beirlant et al. (2004) and de Haan and Ferreira (2007) for textbook treatments of multivariate extreme value theory.
The focus in this paper is on part (ii), which concerns the tail dependence of $\bm{X}$, that is, the probabilities of high values occurring jointly among the variables $X_{j}$. To describe the mathematical theory behind tail dependence, let $C$ be the survival copula of $\bm{X}$: under the standing assumption that the marginal distribution functions $F_{1},\ldots,F_{d}$ are continuous, $C$ is the distribution function of the random vector $\bm{U}=(U_{1},\ldots,U_{d})$ with uniformly distributed components $U_{j}=1-F_{j}(X_{j})$ for $j\in[d]$. Our interest is in high values of $X_{j}$ and thus in low outcomes of $U_{j}$.
@@
The term ‘tail copula’ stems from Schmidt and Stadtmüller (2006) in the bivariate case; Joe et al. (2010) call it ‘tail dependence function’. The tail copula measure already appears in Einmahl et al. (2001), where it is denoted by $\Lambda$. It is closely linked to the exponent measure $\nu$ in Eq. (2.9) below, introduced in de Haan and Resnick (1977) and which constitutes a fundamental notion in multivariate extreme value theory.
@@
Tail copula densities
Throughout, we assume that the tail copula measure $R$ has no mass on the hyperplanes through infinity, that is, $R\left({\left\{{\bm{x}\in\mathbb{E}:x_{j}=\infty}\right\}}\right)=0$ for all $j\in[d]$, so that $R$ is concentrated on $(0,\infty)^{d}$. Furthermore, we assume that $R$ is absolutely continuous with respect to the $d$-dimensional Lebesgue measure with continuous density $r:(0,\infty)^{d}\to[0,\infty)$, that is, $R(B)=\int_{B}r(\bm{x})\,\mathrm{d}\bm{x}$ for all Borel sets $B\subseteq(0,\infty)^{d}$. Choosing $B=\left(\bm{0},\bm{x}\right]$ for $\bm{x}\in(0,\infty)^{d}$, we recover $r$ from $R$ by $r(\bm{x})=\frac{\partial^{d}}{\partial x_{1}\cdots\partial x_{d}}R(\bm{x})$ for all $\bm{x}\in(0,\infty)^{d}$. The marginal constraint (2.2) implies
@@
Properties (2.4) and (2.5) characterise the set of $d$-variate tail copula densities; see Section 4 for some well-known parametric families.
By Scheffé’s lemma from measure theory, if the copula $C$ has density $c$ and the tail copula measure $R$ has density $r$, then (2.1) is implied by $\lim_{t\searrow 0}t^{1-d}c(t\bm{x})=r(\bm{x})$ for all $\bm{x}\in(0,\infty)^{d}$. A word of caution: for a given copula density $c$, the limit $\lim_{t\searrow 0}t^{1-d}c(t\bm{x})$ may exist for all $\bm{x}\in(0,\infty)^{d}$ but not be a tail copula density, as the marginal constraint (2.4) may fail; a case in point is the independence copula, with density $c\equiv 1$, in which case the said limit is zero.
Margins of tail copulas
For a non-empty set $J\subseteq[d]$, let $\pi_{J}:(0,\infty)^{d}\to(0,\infty)^{J}$ denote the coordinate projection $\bm{x}\mapsto\pi_{J}(\bm{x})=\bm{x}_{J}$ and let $R_{J}=R\circ\pi_{J}^{-1}$ denote the $J$-th marginal measure of $R$,
@@
for Borel sets $B\subseteq(0,\infty)^{J}$. The choice $B=\prod_{j\in J}\left(0,x_{j}\right]$ for $x_{j}\in(0,\infty)$ shows that $R_{J}$ is the tail copula measure of the copula $C_{J}$ of $\bm{U}_{J}=\pi_{J}(\bm{U})$, as in Definition 2.1. For $\bm{x}_{J}\in(0,\infty]^{J}\setminus\left\{{\bm{\infty}}\right\}$, we have
$R_{J}(\bm{x}_{J})=R(\tilde{\bm{x}}_{J})\quad\text{with}\quad\tilde{x}_{j}=\begin{dcases}x_{j}&\text{if $j\in J$,}\\ \infty&\text{if $j\in[d]\setminus J$.}\end{dcases}$
The density $r_{J}$ of the measure $R_{J}$ is
@@
for $\bm{x}_{J}\in(0,\infty)^{J}$ and, provided $J\neq[d]$, is obtained from $r$ by integrating out the variables $x_{j}$ with indices $j\not\in J$.
Relations between tail copulas and other concepts in multivariate extremes
For later reference, we review a number of related terms from multivariate extreme value theory. This paragraph can be skipped at first reading, as the contributions in Sections 3 and 5 do not depend on it.
@@
$\Lambda(\,\cdot\,)=R\left({\left\{{\bm{x}\in\mathbb{E}:(1/x_{1},\ldots,1/x_{d})\in\cdot\,}\right\}}\right),$ (2.9)
on $[0,\infty)^{d}\setminus\left\{{\bm{0}}\right\}$, that is, $t\operatorname{\mathsf{P}}(\bm{V}\in tB)\to\Lambda(B)$ as $t\to\infty$ for Borel sets $B$ contained in $[0,\infty)^{d}\setminus[0,\varepsilon]^{d}$ for some $\varepsilon>0$ and satisfying $\Lambda(\partial B)=0$, with $\partial B$ the boundary of $B$. The measure $\Lambda$ is called exponent measure because of the identity $G(\bm{z})=\exp\left\{{-\Lambda\left({\left[\bm{0},\bm{\infty}\right)\setminus\left[{\bm{0},\bm{z}}\right]}\right)}\right\}$ for $\bm{z}\in(0,\infty)^{d}$.
Another statement equivalent to (2.1) is that the conditional distribution of $\bm{V}/t$ given $\max\bm{V}>t$ is asymptotically multivariate Pareto for high threshold $t$, that is, we have the weak convergence relation
@@
Up to a location shift, multivariate Pareto distributions are a special case of multivariate generalised Pareto distributions introduced in Rootzén and Tajvidi (2006). They play a prominent role in the theory of graphical models for extremes (Engelke and Hitz, 2020). In the context of tail copula measures, it is more convenient to take component-wise reciprocals in (2.10) leading to inverted multivariate Pareto distributions, introduced in Section 6.
If the tail copula measure $R$ is concentrated on $(0,\infty)^{d}$ and has density $r$, the exponent measure and the multivariate Pareto distribution have densities too. The exponent measure $\nu$ is concentrated on $(0,\infty)^{d}$ and has density
@@
By (2.3) and (2.9), the exponent measure $\nu$ is homogeneous of order $-1$ and thus determined by the angular measure $H$ on the unit simplex $\Delta_{d-1}=\left\{{\bm{x}\in[0,1]^{d}:\lVert{\bm{x}}\rVert_{1}=1}\right\}$ via
for $t>0$ and Borel sets $B\subseteq\Delta_{d-1}$, with $\lVert{\bm{x}}\rVert_{1}=\sum_{j=1}^{d}|x_{j}|$ for $\bm{x}\in\mathbb{R}^{d}$. Under the above smoothness assumption, the angular measure $H$ is concentrated on the relative interior of the unit simplex $\Delta_{d-1}$ and its density $h$ satisfies $h(\bm{w})=\lambda(\bm{w})$ for $\bm{w}\in\Delta_{d-1}$ (Dombry et al., 2016), see also Lemma B.1 in Appendix B.
@@ 2.2 Regular vine sequences and the vine telescoping product formula
Informally, a regular vine sequence on $d$ elements consists of a linked set of $d-1$ trees, where the edges in tree $j-1$ become nodes in tree $j$, joined by an edge only if they share a common node as edges in tree $j-1$, for $j\in\left\{{2,\ldots,d-1}\right\}$ (Bedford and Cooke, 2002). Regular vine copulas, also called pair copula constructions, allow for flexible high-dimensional modelling using parametric families of bivariate copulas as building blocks for the edges of each tree. For recent overviews, see, for example, the monograph by Czado (2019) or the survey article by Czado and Nagler (2022). Our objective is to do the same for tail copula densities.
Recall that a tree $T=(N,E)$ is a connected, acyclic graph comprising a node set $N$ and an edge set $E\subseteq\left\{{\left\{{x,y}\right\}:\;x,y\in N,\,x\neq y}\right\}$. Any two distinct nodes in a tree are connected by a unique path.
Definition 2.2 (Regular vine tree sequence).
@@
For brevity, we also use the term regular vine sequence. The proximity condition and other concepts defined below are illustrated via a five-dimensional example in Appendix A.
@@
Let $\mathcal{V}=(T_{1},\ldots,T_{d-1})$ be a regular vine sequence. For any edge $e\in E_{j}$, the complete union of $e$ is $A_{e}=e$ if $j=1$ and $A_{e}=A_{a}\cup A_{b}$ if $e=\{a,b\}$ and $j\in\left\{{2,\ldots,d-1}\right\}$; informally, $A_{e}$ is the subset of nodes in $[d]$ reachable from $e$ by the membership relation. The conditioning set of an edge $e=\{a,b\}\in E_{j}$ with $j\geqslant 2$ is $D_{e}=A_{a}\cap A_{b}$ and the conditioned set of $e$ is $\mathcal{C}_{e}=\mathcal{C}_{e,a}\cup\mathcal{C}_{e,b}$ with $\mathcal{C}_{e,a}=A_{a}\setminus D_{e}=A_{a}\setminus A_{b}$ and $\mathcal{C}_{e,b}=A_{b}\setminus D_{e}=A_{b}\setminus A_{a}$; if $e=\{a,b\}\in E_{1}$, we have $\mathcal{C}_{e,a}=\{a\}$, $\mathcal{C}_{e,b}=\{b\}$, $\mathcal{C}_{e}=e$ and $D_{e}=\varnothing$.
Clearly, $D_{e}=A_{a}\setminus\mathcal{C}_{e}=A_{b}\setminus\mathcal{C}_{e}$. Further, $\mathcal{C}_{e,a}$, $\mathcal{C}_{e,b}$ and $D_{e}$ are disjoint subsets of $N_{1}$; $\mathcal{C}_{e,a}$ and $\mathcal{C}_{e,b}$ are singletons while $D_{e}$ has $j-1$ elements; and $\mathcal{C}_{e,a}\cup\mathcal{C}_{e,b}\cup D_{e}=A_{e}$, a node set with $j+1$ elements. If $j\geqslant 2$, then $D_{a}\cup D_{b}\subseteq D_{e}$; indeed, by proximity, $a\cap b=\left\{{k}\right\}$ with $k\in N_{j-1}$, and $D_{a}\subset A_{k}\subset A_{a}\cap A_{b}=D_{e}$; similarly $D_{b}\subset D_{e}$. Further, the conditioned sets are related by $\mathcal{C}_{e,a}=\mathcal{C}_{e}\cap\mathcal{C}_{a}$ and $\mathcal{C}_{e,b}=\mathcal{C}_{e}\cap\mathcal{C}_{b}$. To see this, note that, by proximity, we have $a=\left\{{a_{1},k}\right\}$ and $b=\left\{{b_{1},k}\right\}$ for three distinct elements $a_{1},b_{1},k\in N_{j-1}$, and $\mathcal{C}_{e,a}=A_{a}\setminus A_{b}=\left({A_{a_{1}}\cup A_{k}}\right)\setminus\left({A_{b_{1}}\cup A_{k}}\right)\subseteq A_{a_{1}}\setminus A_{k}=\mathcal{C}_{a,a_{1}}$. Since both the left- and right-hand side are singletons, we must have $\mathcal{C}_{e,a}=\mathcal{C}_{a,a_{1}}=\mathcal{C}_{e}\cap\mathcal{C}_{a}$; similarly $\mathcal{C}_{e,b}=\mathcal{C}_{b,b_{1}}=\mathcal{C}_{e}\cap\mathcal{C}_{b}$.
The edge $e$ is also written as $e=\left({\mathcal{C}_{e,a},\mathcal{C}_{e,b};D_{e}}\right)$ and, since $\mathcal{C}_{e,a}$ and $\mathcal{C}_{e,b}$ are singletons, $\{a_{e}\}$ and $\{b_{e}\}$, say, we abbreviate $e=\left({a_{e},b_{e};D_{e}}\right)$, with $a_{e},b_{e}\in N_{1}$ and the convention that $a_{e}<b_{e}$, that is, $a_{e}=\min\mathcal{C}_{e}$ and $b_{e}=\max\mathcal{C}_{e}$. Similarly, for $e=\left\{{a,b}\right\}\in E_{1}$, the convention is that $a<b$. In Fig. 1, for instance, the edge $e$ labelled $(14;23)$ in $T_{3}$ has $a_{e}=1$, $b_{e}=4$, and $D_{e}=\left\{{2,3}\right\}$. Appendix A contains a fully worked out example to illustrate these concepts.
For positive scalars $\gamma_{1},\ldots,\gamma_{d}$ with $\gamma_{1}=1$, the identity $\prod_{j=2}^{d}\gamma_{j}/\gamma_{j-1}=\gamma_{d}$ is known as a telescoping product. Regular vine sequences enjoy a similar property that will be key to Theorem 5.1 and which, to the best of our knowledge, has not yet been stated in the literature.
@@
Let $d\geqslant 3$ be an integer and let $\gamma_{J}\in(0,\infty)$ for every non-empty $J\subseteq[d]$ with $\gamma_{j}:=\gamma_{\{j\}}=1$ for every $j\in[d]$. Given a regular vine sequence $\mathcal{V}=\left({T_{j}=\left({N_{j},E_{j}}\right)}\right)_{j=1}^{d-1}$ on $[d]$, we have
@@
The formula (2.13) is spelled out in Eq. (A.1) in Appendix A for a five-dimensional regular vine sequence.
Remark 2.5 (Edge-wise vine telescoping product).
$\gamma_{A_{f}}=\prod_{e\in E_{1}:A_{e}\subseteq A_{f}}\gamma_{e}\cdot\prod_{i=2}^{j}\prod_{e\in E_{i}:A_{e}\subseteq A_{f}}\frac{\gamma_{\{a_{e},b_{e}\}|D_{e}}}{\gamma_{a_{e}|D_{e}}\cdot\gamma_{b_{e}|D_{e}}}.$ (2.15)

### v2

@@ 1 Introduction
For copula-based dependence modelling, outside the context of extreme value analysis, vine constructions have grown into a versatile and widely applied approach (Czado, 2019; Czado and Nagler, 2022; Nagler and Vatter, 2023). Our contribution is to make vine-based methods operational, in theory and practice, for densities of exponent measures of multivariate extreme value distributions. The challenge to overcome is that these densities do not integrate to one but to infinity. A change of margin transforms these into tail copula densities, which still have infinite mass, but with structural properties that resemble copula densities more closely.
@@
The models constructed in this way turn out to be the tail limits of regular vine copulas. For the special case of D-vine copulas, such tail limits were already computed in Joe (1996) and Joe et al. (2010), focusing on tail dependence functions rather than densities. Li and Wu (2013) define tail densities and compute those of D-vine copulas in dimensions three and four. In Simpson et al. (2021), tails of D-vine and C-vine copulas were investigated from the perspective of the limiting shapes of sample clouds.
X-vine constructions are related to but different from recently proposed graphical models for extremes, either for directed or undirected graphs (Gissibl and Klüppelberg, 2018; Engelke and Hitz, 2020; Engelke et al., 2022; Engelke et al., 2024). Markov trees (Segers, 2020; Hu et al., 2024) are a special case in the intersection of graphical and X-vine models.
After reviewing background on multivariate extremes and regular vine sequences in Section 2, we state a version of Sklar’s theorem for tail copula densities in Section 3. Some parametric examples are worked out in Section 4, with a focus on the simplifying assumption that the conditional copula densities only depend on the indices of the conditioning variables, but not their actual values. Section 5 contains the paper’s main results, showing on the one hand how a general tail copula density can be decomposed along any regular vine sequence, and, on the other hand, how to construct a tail copula density from a regular vine sequence and bivariate building blocks. The so-called X-vine models arising in this way are put to work in the subsequent sections, covering simulation algorithms (Section 6), parameter estimation and model selection (Section 7), simulation studies (Section 8), and a case study on US flight delay data (Section 9) taken from Hentschel et al. (2022). Section 10 concludes. The supplementary material contains a detailed example to illustrate our theory, the proofs of the paper’s results, expressions based on exponent measure densities, and additional numerical results. The methods are implemented in the R (R Core Team, 2023) package Xvine [^1] Available from https://github.com/JeongjinLee88/Xvine/tree/main, relying in particular on packages graphicalExtremes (Engelke et al., 2022) and VineCopula (Nagler et al., 2023).
@@ 2.1 Multivariate extreme value theory: tail copulas and their densities
Tail copulas.
Classical extreme value theory starts from the assumption that the distribution function $F$ of a random vector $\bm{X}=(X_{1},\ldots,X_{d})$ is in the max-domain of attraction of a multivariate extreme value distribution (Beirlant et al., 2004; de Haan and Ferreira, 2007). This assumption concerns the tails of the univariate marginal distribution functions $F_{1},\ldots,F_{d}$ and the tail dependence structure of $\bm{X}$. Our focus is on the latter aspect, that is, on probabilities of high values occurring jointly among the variables $X_{j}$. Let $C$ be the survival copula of $\bm{X}$: under the standing assumption that $F_{1},\ldots,F_{d}$ are continuous, $C$ is the distribution function of the random vector $\bm{U}=(U_{1},\ldots,U_{d})$ with uniformly distributed components $U_{j}=1-F_{j}(X_{j})$ for $j\in[d]$. Our interest is in high values of $X_{j}$ and thus in low outcomes of $U_{j}$.
@@
The term ‘tail copula’ stems from Schmidt and Stadtmüller (2006) in case $d=2$. The tail copula measure already appears in Einmahl et al. (2001) and is closely linked to the exponent measure $\Lambda$ in Eq. (2.9) below, introduced in de Haan and Resnick (1977). The limit (2.1) appears in Jaworski (2006) and in Joe et al. (2010), who call it ‘tail dependence function’.
@@
Any Borel measure $R$ on $\mathbb{E}$ satisfying (2.2) and (2.3) is a tail copula measure, i.e., is the limit in (2.1) for some copula $C$; see Lemma B.1 in the supplement.
Tail copula densities.
Throughout, we assume that the tail copula measure $R$ has no mass on the hyperplanes through infinity, $R\left({\left\{{\bm{x}\in\mathbb{E}:x_{j}=\infty}\right\}}\right)=0$ for all $j\in[d]$, so $R$ is supported on $(0,\infty)^{d}$. Further, we assume that $R$ is absolutely continuous with respect to the $d$-dimensional Lebesgue measure with continuous density $r:(0,\infty)^{d}\to[0,\infty)$, that is, $R(B)=\int_{B}r(\bm{x})\,\mathrm{d}\bm{x}$ for all Borel sets $B\subseteq(0,\infty)^{d}$. Choosing $B=\left(\bm{0},\bm{x}\right]$ for $\bm{x}\in(0,\infty)^{d}$, we recover $r$ from $R$ by $r(\bm{x})=\frac{\partial^{d}}{\partial x_{1}\cdots\partial x_{d}}R(\bm{x})$ for all $\bm{x}\in(0,\infty)^{d}$. The marginal constraint (2.2) implies
@@
Properties (2.4) and (2.5) characterise the set of $d$-variate tail copula densities; see Section 4 for some parametric families.
By Scheffé’s lemma, if the copula $C$ has density $c$ and the tail copula measure $R$ has density $r$, then (2.1) is implied by $\lim_{t\searrow 0}t^{1-d}c(t\bm{x})=r(\bm{x})$ for all $\bm{x}\in(0,\infty)^{d}$. Li and Wu (2013) call $r$ the (lower) tail density function of $C$. A word of caution: for a given copula density $c$, the limit $\lim_{t\searrow 0}t^{1-d}c(t\bm{x})$ may exist for all $\bm{x}\in(0,\infty)^{d}$ but not be a tail copula density, as the marginal constraint (2.4) may fail. A case in point is the independence copula, with density $c\equiv 1$, in which case the said limit is zero.
Margins of tail copulas.
For non-empty $J\subseteq[d]$, let $\pi_{J}:(0,\infty)^{d}\to(0,\infty)^{J}$ denote the coordinate projection $\bm{x}\mapsto\pi_{J}(\bm{x})=\bm{x}_{J}$ and let $R_{J}=R\circ\pi_{J}^{-1}$ denote the $J$-th marginal measure
@@
for Borel sets $B\subseteq(0,\infty)^{J}$. The choice $B=\prod_{j\in J}\left(0,x_{j}\right]$ for $x_{j}\in(0,\infty)$ shows that $R_{J}$ is the tail copula measure of the copula $C_{J}$ of $\bm{U}_{J}=\pi_{J}(\bm{U})$, as in Definition 2.1. For $\bm{x}_{J}\in(0,\infty]^{J}\setminus\left\{{\bm{\infty}}\right\}$, we have $R_{J}(\bm{x}_{J})=R(\tilde{\bm{x}}_{J})$ with $\tilde{x}_{j}=x_{j}$ if $j\in J$ and $\tilde{x}_{j}=\infty$ if $j\in[d]\setminus J$. For $\bm{x}_{J}\in(0,\infty)^{J}$, the density $r_{J}$ of the measure $R_{J}$ is
@@
and, provided $J\neq[d]$, is obtained from $r$ by integrating out the variables $x_{j}$ with indices $j\not\in J$.
Relations between tail copulas and other concepts in multivariate extremes.
We review a number of related terms from multivariate extreme value theory. This paragraph can be skipped at first reading, as the contributions in Sections 3 to 5 do not depend on it.
@@
$\Lambda(\,\cdot\,)=R\left({\left\{{\bm{x}\in\mathbb{E}:(1/x_{1},\ldots,1/x_{d})\in\cdot\,}\right\}}\right)\qquad\text{ on }[0,\infty)^{d}\setminus\left\{{\bm{0}}\right\},$ (2.9)
that is, $t\operatorname{\mathsf{P}}(\bm{V}\in tB)\to\Lambda(B)$ as $t\to\infty$ for Borel sets $B$ contained in $[0,\infty)^{d}\setminus[0,\varepsilon]^{d}$ for some $\varepsilon>0$ and satisfying $\Lambda(\partial B)=0$, with $\partial B$ the boundary of $B$. The measure $\Lambda$ is called exponent measure because of the identity $G(\bm{z})=\exp\left\{{-\Lambda\left({\left[\bm{0},\bm{\infty}\right)\setminus\left[{\bm{0},\bm{z}}\right]}\right)}\right\}$ for $\bm{z}\in(0,\infty)^{d}$.
Another statement equivalent to (2.1) is that the conditional distribution of $\bm{V}/t$ given $\max\bm{V}>t$ is asymptotically multivariate Pareto that is, we have the weak convergence
@@
Up to a location shift, multivariate Pareto distributions are a special case of multivariate generalised Pareto distributions introduced in Rootzén and Tajvidi (2006). They play a prominent role in the theory of graphical models for extremes (Engelke and Hitz, 2020).
In practice, it may be convenient to reduce the information contained in the tail copula measure to a dependence coefficient. In this paper, we will encounter the tail dependence coefficients $\chi_{J}=R(\{\bm{x}\in(0,\infty]^{d}:\max_{j\in J}x_{j}<1\})=R_{J}(1,\ldots,1)$ for index sets $J\subseteq[d]$ with two or more elements.
If the tail copula measure $R$ is concentrated on $(0,\infty)^{d}$ and has density $r$, the exponent measure and the multivariate Pareto distribution have densities too. The exponent measure $\Lambda$ is concentrated on $(0,\infty)^{d}$ and has density
@@ 2.2 Regular vine sequences and the vine telescoping product formula
Informally, a regular vine sequence on $d$ elements consists of a linked set of $d-1$ trees, where the edges in tree $j-1$ become nodes in tree $j$, joined by an edge only if they share a common node as edges in tree $j-1$, for $j\in\left\{{2,\ldots,d-1}\right\}$ (Bedford and Cooke, 2002). Regular vine copulas, also called pair copula constructions, allow for flexible high-dimensional modelling using parametric families of bivariate copulas as building blocks for the edges of each tree. For recent overviews, see, for example, Czado (2019) or Czado and Nagler (2022). Our objective is to do the same for tail copula densities. All concepts introduced below are illustrated via a five-dimensional example in Section A of the supplementary material.
A tree $T=(N,E)$ is a connected, acyclic graph comprising a finite node set $N$ and an edge set $E\subseteq\left\{{\left\{{x,y}\right\}:\;x,y\in N,\,x\neq y}\right\}$. Any two distinct nodes in a tree are connected by a unique path.
Definition 2.2 (Regular vine (tree) sequence).
@@
Let $\mathcal{V}=(T_{1},\ldots,T_{d-1})$ be a regular vine sequence, with $T_{j}=(N_{j},E_{j})$. For any edge $e\in E_{j}$, the complete union of $e$ is $A_{e}=e$ if $j=1$ and $A_{e}=A_{a}\cup A_{b}$ if $e=\{a,b\}$ and $j\in\left\{{2,\ldots,d-1}\right\}$. Informally, $A_{e}$ is the subset of nodes in $[d]$ reachable from $e$ by the membership relation. The conditioning set of an edge $e=\{a,b\}\in E_{j}$ with $j\geqslant 2$ is $D_{e}=A_{a}\cap A_{b}$ and the conditioned set of $e$ is $\mathcal{C}_{e}=\mathcal{C}_{e,a}\cup\mathcal{C}_{e,b}$ with $\mathcal{C}_{e,a}=A_{a}\setminus D_{e}=A_{a}\setminus A_{b}$ and $\mathcal{C}_{e,b}=A_{b}\setminus D_{e}=A_{b}\setminus A_{a}$. If $e=\{a,b\}\in E_{1}$, then $\mathcal{C}_{e,a}=\{a\}$, $\mathcal{C}_{e,b}=\{b\}$, $\mathcal{C}_{e}=e$ and $D_{e}=\varnothing$.
For an edge $e=\left\{{a,b}\right\}\in E_{j}$, the sets $\mathcal{C}_{e,a}$, $\mathcal{C}_{e,b}$ and $D_{e}$ are disjoint and their union equals $A_{e}$, a node set with $j+1$ elements. The sets $\mathcal{C}_{e,a}$ and $\mathcal{C}_{e,b}$ are singletons while $D_{e}$ has $j-1$ elements. Clearly, $D_{e}=A_{a}\setminus\mathcal{C}_{e}=A_{b}\setminus\mathcal{C}_{e}$. The edge $e$ is also written as $e=\left({\mathcal{C}_{e,a},\mathcal{C}_{e,b};D_{e}}\right)$ and, since $\mathcal{C}_{e,a}$ and $\mathcal{C}_{e,b}$ are singletons, $\{a_{e}\}$ and $\{b_{e}\}$, say, we abbreviate $e=\left({a_{e},b_{e};D_{e}}\right)$, with $a_{e},b_{e}\in[d]$ and the convention that $a_{e}<b_{e}$, that is, $a_{e}=\min\mathcal{C}_{e}$ and $b_{e}=\max\mathcal{C}_{e}$. Similarly, for $e=\left\{{a,b}\right\}\in E_{1}$, the convention is that $a<b$. In Fig. 1, for instance, the edge $e$ labelled $(14;23)$ in $T_{3}$ has $a_{e}=1$, $b_{e}=4$, and $D_{e}=\left\{{2,3}\right\}$. See Section B.1 in the supplement for some additional properties.
For positive scalars $\gamma_{1},\ldots,\gamma_{d}$, the identity $\prod_{j=2}^{d}(\gamma_{j}/\gamma_{j-1})=\gamma_{d}/\gamma_{1}$ is a telescoping product. Regular vine sequences enjoy a similar property that will be key to Theorem 5.1 and which, to the best of our knowledge, has not yet been stated in the literature.
@@
Let $d\geqslant 3$ be an integer and let $\gamma_{J}\in(0,\infty)$ for every non-empty $J\subseteq[d]$, with $\gamma_{j}:=\gamma_{\{j\}}=1$ for every $j\in[d]$. Given a regular vine sequence $\mathcal{V}=\left({T_{j}=\left({N_{j},E_{j}}\right)}\right)_{j=1}^{d-1}$ on $[d]$, we have
@@
Remark B.4 in Appendix B.1 in the supplement generalises Lemma 2.4 to factorisations of $\gamma_{J}$ for certain non-empty subsets $J\subseteq[d]$, while Remark B.5 provides a generalisation of (2.13) without the restriction that $\gamma_{j}=1$ for all $j\in[d]$.